9. In the past few years, a few new "private garage and track" facilities have opened in various places across the U.S. For example, in Michigan, the M1 Concourse (Links to an external site.) facility opened recently opened.

At the basic level, these facilities charge consumers an annual membership fee and allow them to have a dedicated garage space to store their car. (Usually a hot rod or track car like a Corvette, Porsche, or Ferrari.) Members then also have access to a track where they can race their car.

Suppose that you are the pricing director at one of these facilities and you're trying to figure out how much to charge, both for the annual membership fee and for each hour of track time that a member uses.

Suppose you know that the market of potential members can be divided into two groups: die-hards and enthusiasts. Each "die-hard" consumer has an annual demand for track time that can be given by Q = 130 - P. Each "enthusiast" consumer has an annual demand for track time that can be given by Q = 100 - P. In both equations P is the price per hour of track time and Q is the number of hours of track time demanded per year.

The facility has fixed costs, and the marginal cost of each hour of track time used is only $10.

Now suppose that you can't tell the customer groups apart. That is, you don't know what "type" any potential member is and you have no way of charging them different prices. Suppose that you decide to set the per-hour track price at your marginal cost. Finally, suppose that you know that 50% of your potential customers are "die-hards" and 50% are "enthusiasts".

  • To maximize profits, what should you set the annual membership fee to?
    (a) $ ________________ per year
  • What would you do if, instead, you knew that 60% of your potential customers are "die-hards" and only 40% are "enthusiasts"?
  • To maximize profits, what should you set the annual membership fee to?
    (b) $ ______________ per year
Answer #1

a) To maximise profits Q+P=100+130=230 UNITS

=230*40/100=230*.0.04=9.2 UNITS Per year

A firm maximizes profit by operating where marginal revenue equals marginal cost. In the short run, a change in fixed costs has no effect on the profit maximizing output or price. The firm merely treats short term fixed costs as sunk costs and continues to operate as before.There are several perspectives one can take on this problem. First, since profit equals revenue minus cost, one can plot graphically each of the variables revenue and cost as functions of the level of output and find the output level that maximizes the difference (or this can be done with a table of values instead of a graph). Second, if specific functional forms are known for revenue and cost in terms of output, one can use calculus to maximize profit with respect to the output level.

since 60% of the customers are die hard=60/100*130=0.06*130=7.8 units per year

A Regular Member pays an annual membership fee determined by the Executive Committee.

We have two membership rates, a normal rate and a student rate.

Criterion for student rate is being enrolled in a regular program of study related to the subjects of ITEA, with a maximum of 3 years at this rate.

The membership fees for the year 2017-2018 are:

  • EUR 70 for normal rate
  • EUR 40 for student rate

Membership during the year 2017 – 2018 will entitle member to:

  • a discount of EUR 70 / 40 on registration for the Annual ITEA Conference and/or School in Barcelona, June 19 – 23, 2017
  • a discount of EUR 70 / 40 on registration for the Annual ITEA Conference and/or School in Hong Kong, June 25 – 29, 2018
  • online access to ECOTRA (Economics of Transportation: Journal of the ITEA)